OpenStudy (anonymous):

If f(x) = 3x + 1, then f(a + h) - f(a) = how do I do these problems??

OpenStudy (amistre64):

how would you do: f(3) - f(8) ??

OpenStudy (amistre64):

do you even know how to define something like f(3) ??

OpenStudy (hba):

You just have to substitute in place of 'x'

OpenStudy (anonymous):

so I'd substiture the 3x+1 into what?

OpenStudy (amistre64):

your not dubbing in the entire function, your subbing the x part of the function by the definition of x

OpenStudy (amistre64):

*subbing in

OpenStudy (amistre64):

spose we are given some function: g(x) ; what would this look like when x=2? replace x by 2 to get: g(2) now spose we know the rule for g(x) is x+4; then g(2) is the value of the function when x=2; 2+4 = 6

OpenStudy (hba):

Suppose*

OpenStudy (amistre64):

in your setup, they ask you to sub in x=a and x=a+h, then subtract one from the other

OpenStudy (hba):

^ Exactly.

OpenStudy (amistre64):

does this make sense?

OpenStudy (anonymous):

Yes it's starting to thank you

OpenStudy (amistre64):

good, then lets try to work it thru with you; lets define f(a) , which tells us that x=a. what do we get?

OpenStudy (amistre64):

f(x) = 3x + 1 , i like to circle the x just to keep track of it f(x) = 3(x) + 1 ; now, when x=a we get f(a) = 3(a) + 1 f(x) = 3(x) + 1 ; now, when x=(a+h) we get f(a+h) = 3(a+h) + 1 so, what is f(a+h)-f(a)? ( 3(a+h) + 1 ) - ( 3(a) + 1 ) and simplify

OpenStudy (hba):

@amistre64 I do not think he knows what going in his class though.

OpenStudy (amistre64):

thats fine, knowing takes time and practice :)

OpenStudy (anonymous):

Thank you!

OpenStudy (hba):

I guess so.

OpenStudy (amistre64):

youre welcome, and good luck with it ;)